# On the Martin boundary for discrete TASEP

### v1

@@ 3.3. Independence of $M^{\Phi}_{n}$
De Finetti’s theorem implies that the set of extreme points of $\mathcal{M}_{sym}(\mathbb{Z}_{\geq 0}^{\infty})$ consists of the product measures, see [HS55, Theorem 5.3].
@@ 3.4. Examples of $M^{\Phi}_{n}$
Assuming that $s=0$ and $\Phi(0)\neq 0$, we can use Propositions 3.9, 3.11 to get
@@
When $p=0$ the first three examples degenerate to extreme coherent systems on the positive part of Gelfand-Tsetlin graph and they fully describe its boundary, see [Ols03]. In this case the coherent systems are related to characters of the infinite unitary group, and the name for the third example comes from the connection to Plancherel measures on the unitary group. However, the coherent system from the last example does not have any analogue in the $p=0$ case: $\Phi(z)=\frac{1-p}{1-pz}$ degenerates to $1$ when $p=0$.
@@ 4.1. Integral formulae for $G^{(0,\mathfrak{b})}_{\lambda}$
Our asymptotic analysis is based on two contour integral formulae. The first formula comes from the fact that the functions $G_{\lambda}^{(0,\mathfrak{b})}$ are degenerations of spin $q$-Whittaker functions $\mathbb{F}_{\lambda/\mu}$, introduced in [BW20]. Here we define these functions following the notation from [BK24]. Let $s,\xi,q$ be a triple of complex parameters. For a single variable $x$ and a pair of partitions $\lambda,\mu$ set
@@
Theorem 4.1 ([BW20, Theorem 8.1],[BK24, Theorem 5.13]).
@@
The functions $G_{\lambda}^{(\mathfrak{a},0)}$ can also be obtained as a degeneration of spin Hall-Littlewood functions $\mathsf{G}_{\lambda}(x_{1},\dots,x_{n}\mid\Xi,\mathsf{S})$ from [BP18]. Comparing [BP18, Theorem 4.14] with (2.1) one can show that
@@
where $l(\lambda)\leq n$ and in the left-hand side $\lambda$ is treated as a signature of length $n$. Then the integral formula for $G^{(\mathfrak{a},0)}_{\lambda}$ obtained in Proposition 4.3 can be seen as a degeneration of the integral formula for $\mathsf{G}_{\lambda}(x_{1},\dots,x_{n}\mid\Xi,\mathsf{S})$ from [BP18, Corollary 7.16].
@@ 4.2. $G^{(0,-p)}_{\lambda}(1^{N})$ with $k$ rows
We use a standard steepest descent argument to establish this asymptotic behavior. The idea is to localize the integral to a small neighborhood of a critical point of the integrand, see e.g. [Cop65] or [Erd56] for the basics of the method. Our first step is to rewrite the integral expression from Proposition 4.3 in an exponential form. Fix the analytic branch of $\ln(u)$ on $\mathbb{C}\backslash\mathbb{R}_{\leq 0}$ with real values on $\mathbb{R}_{>0}$ and set
@@ 5.2. Extremality of $M_{n}^{\mathcal{A},\varnothing}$ and $M_{n}^{\varnothing,\mathcal{B}}$
Our proof follows the lines of [Ols03, Proposition 10.8].
@@
Let $\psi$ be a bounded Borel function on $\mathcal{T}$, which we treat as a random variable with respect to a central measure $M$. Define $\psi_{N}=\mathbb{E}_{M}[\psi\mid\mathcal{B}_{-N}]$ and $\psi_{\infty}=\mathbb{E}_{M}[\psi\mid\mathcal{B}_{-\infty}]$ denote the conditional expectations. Then $\psi_{-N}$ form a reverse martingale. By the reverse martingale convergence theorem [Doo12, Theorem XI.15] we get
@@ 6. Open questions and future directions
Note that when either $k=0$ or $l=0$ we can almost prove Conjecture 6.1 using methods of this paper. By [Ols03, Proposition 10.8] the converse to the second part of Proposition 5.3 is true: if $M$ is an extreme central measure then for $M$-almost every path $(t_{N})$ we should have
@@
From Proposition (5.3) and [Ols03, Proposition 10.8], a closely related question is the description of paths $t=(t_{N})_{N}\in\mathcal{T}$ such that the limit $\lim_{N\to\infty}p_{N,n}(t_{N},\lambda)$ exists for every $\lambda\in\mathbb{G}^{p}_{n}$. This is also equivalent to existence of limits

### v2

@@ 3.3. Independence of $M^{\Phi}_{n}$
In other words, $M^{\Phi}_{1}((n))$ are Taylor coefficients of $\Phi$ and the coherent systems $\{M^{\Phi}_{n}\}_{n}$ are different for different $\Phi\in\mathcal{F}$.
Proposition 3.16.
@@
Proof.
[…]
Now assume that $\Phi_{k}\in\mathcal{F}$ is a sequence of functions such that $\Upsilon(\Phi_{k})$ converge to a coherent system $\{M_{n}\}_{n}$. In other words, $\lim_{k\to\infty}M_{n}^{\Phi_{k}}(\lambda)=M_{n}(\lambda)$ for each $n\geq 0,\lambda\in\mathbb{G}_{n}^{p}$. Consider
$\Phi(z)=\sum_{n\geq 0}M_{1}((1))z^{n}.$
[…]
For the rest of the argument let $a_{n,k}=M^{\Phi_{k}}_{1}((n))$ and $a_{n}=M_{1}((n))$ denote the Taylor coefficients of $\Phi_{k}$ and $\Phi$ respectively. Then $\lim_{k\to\infty}a_{n,k}=a_{n}$. For an arbitrary $\varepsilon>0$ fix sufficiently large $N$ such that $\sum_{n>N}a_{n}<\varepsilon$. Then for sufficiently large $k$ we have $\sum_{n=0}^{N}|a_{n}-a_{n,k}|<\varepsilon$, which implies
$\sum_{n>N}a_{N,k}=1-\sum_{n=0}^{N}a_{N,k}<1-\sum_{n=0}^{N}a_{N}+\varepsilon=\sum_{n>N}^{N}a_{N}+\varepsilon<2\varepsilon.$
Hence for sufficiently large $k$ we get $|\Phi(z)-\Phi_{k}(z)|<4\varepsilon$ for all $|z|\leq 1$. This implies the uniform convergence $\Phi_{k}\rightrightarrows\Phi$ on $\overline{\mathbb{D}}$, which in turn leads to $\Phi\in\mathcal{F}$. Since $M_{n}^{\Phi}(\lambda)$ is continuous in $\Phi$, we get $M_{n}^{\Phi}(\lambda)=\lim_{k\to\infty}M_{n}^{\Phi_{k}}(\lambda)=M_{n}(\lambda)$. ∎
@@
De Finetti’s theorem implies that the set of extreme points of $\mathcal{M}_{sym}(\mathbb{Z}_{\geq 0}^{\infty})$ consists of the product measures, see [deFin, Theorem 5.3].
@@ 3.4. Examples of $M^{\Phi}_{n}$
Assuming that $s=0$ and $\Phi(0)\neq 0$, we can use Propositions 3.10, 3.12 to get
@@
When $p=0$ the first three examples degenerate to extreme coherent systems on the positive part of Gelfand-Tsetlin graph and they fully describe its boundary, see [Olsh01]. In this case the coherent systems are related to characters of the infinite unitary group, and the name for the third example comes from the connection to Plancherel measures on the unitary group. However, the coherent system from the last example does not have any analogue in the $p=0$ case: $\Phi(z)=\frac{1-p}{1-pz}$ degenerates to $1$ when $p=0$.
@@ 4.1. Integral formulae for $G^{(0,\mathfrak{b})}_{\lambda}$
Our asymptotic analysis is based on two contour integral formulae. The first formula comes from the fact that the functions $G_{\lambda}^{(0,\mathfrak{b})}$ are degenerations of spin $q$-Whittaker functions $\mathbb{F}_{\lambda/\mu}$, introduced in [BW17]. Here we define these functions following the notation from [BK21]. Let $s,\xi,q$ be a triple of complex parameters. For a single variable $x$ and a pair of partitions $\lambda,\mu$ set
@@
Theorem 4.1 ([BW17, Theorem 8.1],[BK21, Theorem 5.13]).
@@
The functions $G_{\lambda}^{(\mathfrak{a},0)}$ can also be obtained as a degeneration of spin Hall-Littlewood functions $\mathsf{G}_{\lambda}(x_{1},\dots,x_{n}\mid\Xi,\mathsf{S})$ from [BP16]. Comparing [BP16, Theorem 4.14] with (2.1) one can show that
@@
where $l(\lambda)\leq n$ and in the left-hand side $\lambda$ is treated as a signature of length $n$. Then the integral formula for $G^{(\mathfrak{a},0)}_{\lambda}$ obtained in Proposition 4.3 can be seen as a degeneration of the integral formula for $\mathsf{G}_{\lambda}(x_{1},\dots,x_{n}\mid\Xi,\mathsf{S})$ from [BP16, Corollary 7.16].
@@ 4.2. $G^{(0,-p)}_{\lambda}(1^{N})$ with $k$ rows
We use a standard steepest descent argument to establish this asymptotic behavior. The idea is to localize the integral to a small neighborhood of a critical point of the integrand, see e.g. [Copson] or [Erd56] for the basics of the method. Our first step is to rewrite the integral expression from Proposition 4.3 in an exponential form. Fix the analytic branch of $\ln(u)$ on $\mathbb{C}\backslash\mathbb{R}_{\leq 0}$ with real values on $\mathbb{R}_{>0}$ and set
@@ 5.2. Extremality of $M_{n}^{\mathcal{A},\varnothing}$ and $M_{n}^{\varnothing,\mathcal{B}}$
Our proof follows the lines of [Olsh01, Proposition 10.8].
@@
Let $\psi$ be a bounded Borel function on $\mathcal{T}$, which we treat as a random variable with respect to a central measure $M$. Define $\psi_{N}=\mathbb{E}_{M}[\psi\mid\mathcal{B}_{-N}]$ and $\psi_{\infty}=\mathbb{E}_{M}[\psi\mid\mathcal{B}_{-\infty}]$ denote the conditional expectations. Then $\psi_{-N}$ form a reverse martingale. By the reverse martingale convergence theorem [Doob, Theorem XI.15] we get
@@ 6. Open questions and future directions
Note that when either $k=0$ or $l=0$ we can almost prove Conjecture 6.1 using methods of this paper. By [Olsh01, Proposition 10.8] the converse to the second part of Proposition 5.3 is true: if $M$ is an extreme central measure then for $M$-almost every path $(t_{N})$ we should have
@@
From Proposition (5.3) and [Olsh01, Proposition 10.8], a closely related question is the description of paths $t=(t_{N})_{N}\in\mathcal{T}$ such that the limit $\lim_{N\to\infty}p_{N,n}(t_{N},\lambda)$ exists for every $\lambda\in\mathbb{G}^{p}_{n}$. This is also equivalent to existence of limits